# CurbNet: Curb Detection Framework Based on LiDAR Point Cloud Segmentation

### v2

@@ Abstract
Curb detection is a crucial function in intelligent driving, essential for determining drivable areas on the road. However, the complexity of road environments makes curb detection challenging. This paper introduces CurbNet, a novel framework for curb detection utilizing point cloud segmentation. To address the lack of comprehensive curb datasets with 3D annotations, we have developed the 3D-Curb dataset based on SemanticKITTI, currently the largest and most diverse collection of curb point clouds. Recognizing that the primary characteristic of curbs is height variation, our approach leverages spatially rich 3D point clouds for training. To tackle the challenges posed by the uneven distribution of curb features on the xy-plane and their dependence on high-frequency features along the z-axis, we introduce the Multi-Scale and Channel Attention (MSCA) module, a customized solution designed to optimize detection performance. Additionally, we propose an adaptive weighted loss function group specifically formulated to counteract the imbalance in the distribution of curb point clouds relative to other categories. Extensive experiments conducted on 2 major datasets demonstrate that our method surpasses existing benchmarks set by leading curb detection and point cloud segmentation models. Through the post-processing refinement of the detection results, we have significantly reduced noise in curb detection, thereby improving precision by 4.5 points. Similarly, our tolerance experiments also achieved state-of-the-art results. Furthermore, real-world experiments and dataset analyses mutually validate each other, reinforcing CurbNet’s superior detection capability and robust generalizability. The project website is available at: https://github.com/guoyangzhao/CurbNet/.
@@ I Introduction
Research in curb detection varies based on sensor technology, broadly categorized into vision-based and LiDAR-based methods [7]. This includes the use of monocular cameras, stereo vision, and combinations of 2D and 3D LiDAR sensors. Vision-based methods can provide rich contextual information and achieve effective detection results. However, camera performance is greatly influenced by light and weather conditions and fails to provide accurate depth information directly, which is a critical need for autonomous driving applications[8, 9]. Furthermore, the subtle color differences between regular road surfaces and curbs make it difficult for cameras to accurately detect curbs. In contrast, LiDAR sensors demonstrate robustness under various weather and lighting conditions and offer precise distance measurements[10]. Recently, 3D LiDAR has become one of the most important sensors for perceiving 3D environments in autonomous vehicles[11, 12].
@@
Curb detection using LiDAR can be classified into manual feature methods and learning based methods[13]. Manual feature design methods[14, 15] typically analyze geometric relationships such as height and angle changes between adjacent points, given the difference between drivable roads and curbs [16, 17]. Most curb detection methods follow a sequential point cloud processing procedure [18, 19], including stages of candidate region extraction, manual feature setting and clustering, and post-processing for fitting estimation. These methods, due to their interpretability in terms of safety, are widely used in curb detection for autonomous driving systems. However, the reliability of these manually designed rule-based processes is limited in practical applications, as errors in early detection stages can severely impact subsequent recognition performance. Additionally, they heavily rely on manual feature design, necessitating extensive parameter adjustments for various scenarios such as straight roads, curved roads, and different types of intersections, resulting in low practical efficiency and generalization [16, 20].
@@
With the breakthroughs in deep learning applications in perception, its capabilities in automatic feature extraction and learning have significantly reduced the tedium of manual feature design and parameter adjustment[21]. Particularly, in addressing different road scenarios, it has substantially enhanced model recognition performance and robustness[22]. Recently, some researchers have explored curb detection using CNN-based methods, projecting 3D LiDAR point cloud data into 2D images from a Bird’s Eye View (BEV) perspective, followed by processing these images with CNN models to detect curbs[23, 24]. However, this direct projection of 3D LiDAR data can lead to the loss of essential spatial structural information, particularly the crucial height difference information for curb detection[25].
@@
Considering current research in curb detection and the physical properties of 3D LiDAR, we have identified several challenges that need addressing, as shown in Fig. 1: (a) The primary distinguishing feature of curbs is the subtle height variation from the road surface[14], a challenging feature for models to accurately learn. (b) Curb point clouds from LiDAR scanning show significant distribution differences across various distances[26]. (c) Curbs occupy only a small portion of LiDAR point clouds, appearing as long and narrow lines, making effective model training difficult[25].
@@
    Introducing a comprehensive 3D-Curb point cloud dataset based on SemanticKITTI, to our knowledge which is the largest and most diverse currently available.
@@
    Proposing a new Multi-Scale and Channel Attention (MSCA) module and an imbalance loss strategy, tailored to the distribution characteristics of curb point clouds.
@@
    Developing a multi-cluster fitting post-processing approach to further enhance detection performance.
@@
    Achieving state-of-the-art detection results in complex, large-scale intersection scenarios.
@@ II-C Curb Detection Datasets
There is a significant amount of research in LiDAR-based curb detection; however, high-quality datasets with 3D annotations are scarce. Major automotive datasets such as NuScenes [38], KITTI [39], and SemanticKITTI [40] do not include curb annotations. The robustness of deep learning methods is closely related to the volume of data collected under various environmental conditions. These factors have somewhat hindered the application of deep learning methods in this task. [19] created a public dataset for curb detection, comprising 200 scans collected across five different scenes. [23] developed a public dataset for curb detection, including 5200 scans with BEV labels, collected from urban areas. LCDeT[25] introduced a curb dataset for 128-beam LiDAR, containing 6200 frames of point clouds from urban road scenes, both during daytime and nighttime. [26] proposed a method for 3D curb detection and annotation in LiDAR point clouds, effectively reducing manual annotation time by 50%.
The above-mentioned curb point cloud datasets only label the curb category, which cannot be directly used in autonomous driving scenarios, as recognition of other categories such as vehicles and roads is also necessary. Similarly, other categories surrounding the curb, like roads and sidewalks, can also assist the model in more accurately learning curb features. Based on the existing large SemanticKITTI dataset, we added annotations for the curb category, thereby covering a richer array of real road scenes, totaling up to 7100 frames of point cloud data. To our knowledge, this is currently the largest and most comprehensive curb point cloud dataset with annotations relevant to autonomous vehicles (AV). Table. I illustrates the related LiDAR datasets for curb detection.
@@ III Methodology
To the best of our knowledge, this is the largest curb point cloud dataset to date and the only one with 3D annotations. Table. I compares our dataset with other related curb datasets.
@@ III-B Overview of Model Framework
During the feature extraction stage, voxelized features with dimensions $F\times H\times W\times D$ are fed into a 5-layer deep encoder-decoder structure. Each encoder and decoder backbone incorporates the Multi-Scale and Channel Attention (MSCA) module, specifically designed for curb point cloud feature extraction. A stride-2 sparse convolution is employed as the pooling function.
@@ III-C Multi-Scale and Channel Attention (MSCA) Module
Channel Attention. The primary distinguishing feature of curbs is the subtle height difference between the road and the sidewalk, reflected in the z-axis of point cloud data. Unlike general point cloud segmentation algorithms such as Cylinder3D [43], which primarily focus on feature learning in the xy plane, we designed the channel attention module to capture high-frequency features along the z-axis. Initially, a 1×1×D sparse convolution is applied to preliminarily extract channel features. This is followed by an MLP to encode and decode the channel features, enhancing the representation of height features. The output of the MLP is then passed through a softmax function to generate dynamic weights for each channel. These dynamic weights are multiplied with the channel features to produce the final channel attention output.
@@ III-D Loss Group
Here, $\gamma^{i}$ is decomposed into a class-agnostic parameter $\gamma_{a}$ and a class-specific parameter $\gamma_{b}^{i}$. The parameter $\gamma_{a}$ represents the basic focusing factor under balanced data scenarios, while $\gamma_{b}^{i}\geq 0$ is a variable parameter related to the imbalance degree of class $i$. The term $\eta^{i}=\frac{N}{N_{i}}$ where $N$ is the total number of points in the point cloud, and $N_{i}$ is the number of points in class $i$. The value of $g^{i}$ is constrained to the range $[0,1]$, and $1-\eta^{i}$ inversely reflects the weight for low-frequency classes. The hyperparameter $s$ is a scaling factor that determines the upper limit of $\gamma^{i}$.
@@ IV-B Quantitative Results of Curb Detection
In the NRS dataset experiments (refer to Table. II), this study compared classic segmentation algorithms such as PointPillars, U-Net, Swin-Transformer, and CSWin-Transformer, as well as the state-of-the-art curb detection model, LCDeT. Leveraging the specially designed MSCA module for 3D curb scenarios, CurbNet achieved the highest detection performance on the NRS dataset. With the auxiliary training of relevant labels, CurbNet attained Precision, Recall, and F-1 scores of 0.8281, 0.8329, and 0.8308, respectively. Among them, auxiliary training helps improve Precision by 0.5 points, and post-processing helps improve Precision by 1.4 points.
@@
The experiments in the 3D-Curb dataset not only compared classic and advanced deep learning model algorithms but also included three traditional methods of manual feature extraction, as shown in Table. III. Owing to the robust automatic feature extraction capabilities, deep learning methods significantly outperformed in curb detection, surpassing the other methods by over 10 points in Precision, Recall, and F-1 score. Among the deep learning methods, CurbNet surpassed the best-performing supervised learning model on the SemanticKITTI dataset, Cylinder3D, by 2.5 points in Precision, and exceeded the knowledge distillation model PVKD by 1.5 points. Among them, Precision is improved by more than 1 point through auxiliary training, and by 4.5 points through post-processing.
@@
In our study, we conducted Tolerance performance tests on the 3D-Curb dataset, setting four Tolerances ranging from just 0.05m to 0.2m, as shown in Table. IV. Tests were carried out on four models: 3D U-Net, Cylinder 3D, PVKD, and CurbNet (ours). With the increase in error Tolerance, performance metrics improved significantly. At 0.05m Tolerance, Precision improved by an average of 4 points; at 0.1m, by 9 points; at 0.15m, by 12 points; and at 0.2m, by 13 points. As Tolerance increased, performance gains gradually reached saturation. Notably, our CurbNet exceeded 0.95 in the average values of Precision, Recall, and F-1 score at just 0.15m Tolerance. This represents the optimal performance achieved in curb detection based on point cloud segmentation.
@@ IV-C Visualization Results of Curb Detection
Furthermore, we conducted a visual analysis of the results on the NRS dataset, as shown in Fig. 8. This analysis primarily showcases curb detection in five common road scenarios and four special intersection types. Through a comparative visualization of the detection results and ground truth, our method accurately identified the respective curb features.
@@ IV-D Post-Processing Experiment
Figure 9: Parameter adjustment experiments for multiple clustering and fitting. We conduct control variable experiments on the main parameter distance variable Eps and the minimum sample point variable minPts of DBSCAN clustering. The fewer minPts, the better the overall performance.
@@ IV-E Ablation Study
As shown in Table. V, this study conducted comparative ablation experiments focusing on the main loss functions and crucial module designs of our model. We first conducted individual experiments on the employed $\mathcal{L}_{CE}$ (Cross-Entropy Loss), $\mathcal{L}_{FL}$ (Focal Loss), $\mathcal{L}_{ACE}$ (Adaptive Cross-Entropy Loss), and $\mathcal{L}_{IoU}$ (Intersection over Union Loss). Subsequently, we tested combinations of $\mathcal{L}_{CE}$+$\mathcal{L}_{IoU}$ loss, $\mathcal{L}_{FL}$+$\mathcal{L}_{IoU}$ loss and $\mathcal{L}_{ACE}$+$\mathcal{L}_{IoU}$ loss. In the experimental results, due to the design of the $\mathcal{L}_{ACE}$ loss addressing the imbalance in the number of curb point clouds compared to other categories, its disproportionate weight settings caused the model to overly focus on recall during training. However, the interaction with the $\mathcal{L}_{IoU}$ loss led to a more balanced overall performance, achieving optimal detection capabilities. The combination of $\mathcal{L}_{ACE}$+$\mathcal{L}_{IoU}$ loss outperformed the $\mathcal{L}_{CE}$+$\mathcal{L}_{IoU}$ loss group by 1.3 points in Precision and 2.2 points in Recall. It also surpassed the $\mathcal{L}_{FL}$+$\mathcal{L}_{IoU}$ loss group with improvements of 0.7 points in Precision and 2 points in Recall.
Finally, we compared the model’s performance with and without the MSCA module. Under the $\mathcal{L}_{ACE}$+$\mathcal{L}_{IoU}$ loss setting, including the MSCA module significantly enhanced the performance metrics, particularly increasing Recall by 2 points and the F-1 score by 1 point.

### v3

@@ Abstract
Curb detection is a crucial function in intelligent driving, essential for determining drivable areas on the road. However, the complexity of road environments makes curb detection challenging. This paper introduces CurbNet, a novel framework for curb detection utilizing point cloud segmentation. To address the lack of comprehensive curb datasets with 3D annotations, we have developed the 3D-Curb dataset based on SemanticKITTI, currently the largest and most diverse collection of curb point clouds. Recognizing that the primary characteristic of curbs is height variation, our approach leverages spatially rich 3D point clouds for training. To tackle the challenges posed by the uneven distribution of curb features on the xy-plane and their dependence on high-frequency features along the z-axis, we introduce the Multi-Scale and Channel Attention (MSCA) module, a customized solution designed to optimize detection performance. Additionally, we propose an adaptive weighted loss function group specifically formulated to counteract the imbalance in the distribution of curb point clouds relative to other categories. Extensive experiments conducted on 2 major datasets demonstrate that our method surpasses existing benchmarks set by leading curb detection and point cloud segmentation models. Through the post-processing refinement of the detection results, we have significantly reduced noise in curb detection, thereby improving precision by 4.5 points. Similarly, our tolerance experiments also achieve state-of-the-art results. Furthermore, real-world experiments and dataset analyses mutually validate each other, reinforcing CurbNet’s superior detection capability and robust generalizability. The project website is available at: https://github.com/guoyangzhao/CurbNet/.
@@ I Introduction
Research in curb detection varies based on sensor technology, broadly categorized into vision-based and LiDAR-based methods [7]. This includes the use of monocular cameras, stereo vision, and combinations of 2D and 3D LiDAR sensors [8]. Vision-based methods can provide rich contextual information and achieve effective detection results. However, camera performance is greatly influenced by light and weather conditions and fails to provide accurate depth information directly, which is a critical need for autonomous driving applications [9, 10]. Furthermore, the subtle color differences between regular road surfaces and curbs make it difficult for cameras to accurately detect curbs. In contrast, LiDAR sensors demonstrate robustness under various weather and lighting conditions and offer precise distance measurements [11]. Recently, 3D LiDAR has become one of the most important sensors for perceiving 3D environments in autonomous vehicles [12].
@@
Curb detection using LiDAR can be classified into manual feature methods and learning based methods [13]. Manual feature design methods [14, 15] typically analyze geometric relationships such as height and angle changes between adjacent points, given the difference between drivable roads and curbs [16, 17]. Most curb detection methods follow a sequential point cloud processing procedure [18, 19], including stages of candidate region extraction, manual feature setting and clustering, and post-processing for fitting estimation. These methods, due to their interpretability in terms of safety, are widely used in curb detection for autonomous driving systems. However, the reliability of these manually designed rule-based processes is limited in practical applications, as errors in early detection stages can severely impact subsequent recognition performance. Additionally, they rely on manual feature design heavily, necessitating extensive parameter adjustments for various scenarios such as straight roads, curved roads, and different types of intersections, resulting in low practical efficiency and generalization [16, 20].
@@
With the breakthroughs in deep learning applications in perception, its capabilities in automatic feature extraction and learning have significantly reduced the tedium of manual feature design and parameter adjustment [21, 22]. Particularly, in addressing different road scenarios, it has substantially enhanced model recognition performance and robustness [23]. Recently, some researchers have explored curb detection using CNN-based methods, projecting 3D LiDAR point cloud data into 2D images from a Bird’s Eye View (BEV) perspective, followed by processing these images with CNN models to detect curbs [24, 25]. However, this direct projection of 3D LiDAR data can lead to the loss of essential spatial structural information, particularly the crucial height difference information for curb detection [26].
@@
Considering current research in curb detection and the physical properties of 3D LiDAR, we have identified several challenges that need addressing, as shown in Fig.  1: (a) The primary distinguishing feature of curbs is the subtle height variation from the road surface [14], a challenging feature for models to accurately learn. (b) Curb point clouds from LiDAR scanning show significant distribution differences across various distances [27]. (c) Curbs occupy only a small portion of LiDAR point clouds, appearing as long and narrow lines, making effective model training difficult [26].
@@
    We introduce a comprehensive 3D-Curb point cloud dataset based on SemanticKITTI, which is the largest and most diverse currently available to our knowledge.
@@
    We propose a novel Multi-Scale and Channel Attention (MSCA) module and an imbalance loss strategy, tailored to the distribution characteristics of curb point clouds.
@@
    We develop a multi-cluster fitting post-processing approach to further enhance detection performance.
@@
    Our method achieves state-of-the-art detection results in complex, large-scale intersection scenarios.
@@ II-B Deep Learning Methods
Some general point cloud segmentation models [40, 41, 42] have achieved accurate 3D point cloud detection using relatively simple model structures. Among them, 3D U-Net [40] is a popular model for 3D volumetric data processing, adapted for point cloud segmentation by learning the spatial hierarchies of features. Cylinder3D [41] is a voxel-based method specifically designed to handle large-scale point clouds by projecting them into cylindrical grids, enabling more efficient feature extraction in 3D space. PVKD [42] is an improvement of Cylinder3D based on knowledge distillation, focusing on transferring knowledge from a large teacher model to a smaller student network to enhance segmentation performance. However, these models are not specifically designed to extract curb features, which are crucial for curb detection tasks.
@@ II-C Curb Detection Datasets
There is a significant amount of research in LiDAR-based curb detection; however, high-quality datasets with 3D annotations are scarce. Major automotive datasets such as NuScenes [43], KITTI [44], and SemanticKITTI [45] do not include curb annotations. The robustness of deep learning methods is closely related to the volume of data collected under various environmental conditions. These factors have somewhat hindered the application of deep learning methods in this task. [19] created a public dataset for curb detection, comprising 200 scans collected across five different scenes. [24] developed a public dataset for curb detection, including 5200 scans with BEV labels, collected from urban areas. LCDeT [26] introduced a curb dataset for 128-line LiDAR, containing 6200 frames of point clouds from urban road scenes, both during daytime and nighttime. [27] proposed a method for 3D curb detection and annotation in LiDAR point clouds, effectively reducing manual annotation time by 50%.
The above-mentioned curb point cloud datasets only label the curb category, which cannot be directly used in autonomous driving scenarios, as recognition of other categories such as vehicles and roads is also necessary. Similarly, other categories surrounding the curb, like roads and sidewalks, can also assist the model in more accurately learning curb features. Based on the existing large SemanticKITTI dataset, we added annotations for the curb category, thereby covering a richer array of real road scenes, totaling up to 7100 frames of point cloud data. To our knowledge, this is currently the largest and most comprehensive curb point cloud dataset with annotations relevant to autonomous vehicles (AV). Table I illustrates the related LiDAR datasets for curb detection.
@@ III Methodology
To the best of our knowledge, this is the largest curb point cloud dataset to date and the only one with 3D annotations. Table I compares our dataset with other related curb datasets.
@@ III-B Overview of Model Framework
During the feature extraction stage, voxelized features with dimensions $F\times H\times W\times D$ are fed into a 5-layer deep encoder-decoder structure. Each encoder and decoder backbone incorporates the MSCA module, specifically designed for curb point cloud feature extraction. A stride-2 sparse convolution is employed as the pooling function.
@@ III-C Multi-Scale and Channel Attention (MSCA) Module
Formally, given the input feature map $\mathbf{X}\in\mathbb{R}^{F\times H\times W\times D}$, where $F\times H\times W\times D$ denotes the spatial dimensions, we apply convolutions with varying strides to capture multi-scale features. Specifically, the input is processed using sparse convolutions (denoted as SConv) with stride values $s\in\{1,3,5\}$:
$\begin{split}\mathbf{X}_{s1}&=\text{SConv}_{s=1}(\mathbf{X}),\\ \mathbf{X}_{s3}&=\text{SConv}_{s=3}(\mathbf{X}),\\ \mathbf{X}_{s5}&=\text{SConv}_{s=5}(\mathbf{X})\end{split}$ (1)
These convolutions produce feature maps $\mathbf{X}_{s1}$, $\mathbf{X}_{s3}$, and $\mathbf{X}_{s5}$ at different spatial scales. The multi-scale outputs are fused as follows:
$\mathbf{X}_{\text{ms}}=f(\mathbf{X}_{s1},\mathbf{X}_{s3},\mathbf{X}_{s5})$ (2)
where $f(\cdot)$ represents the pyramid fusion operation, typically achieved by concatenating the outputs followed by additional sparse convolutions for deep integration.
Channel Attention. The primary distinguishing feature of curbs is the subtle height difference between the road and the sidewalk, reflected in the z-axis of point cloud data. Unlike general point cloud segmentation algorithms such as Cylinder3D [41], which primarily focus on feature learning in the xy plane, we designed the channel attention module to capture high-frequency features along the z-axis. Initially, a 1×1×D sparse convolution is applied to preliminarily extract channel features:
$\mathbf{C}=\text{SConv}_{1\times 1\times D}(\mathbf{X})$ (3)
The channel features are then processed through an encoder-decoder structured MLP to further refine the height feature:
$\mathbf{C}_{\text{encoded}}=\text{MLP}_{\text{enc}}(\mathbf{C}),\quad\mathbf{C}_{\text{decoded}}=\text{MLP}_{\text{dec}}(\mathbf{C}_{\text{encoded}})$ (4)
The output of the MLP, $\mathbf{C}_{\text{decoded}}$, is then passed through two parallel branches. In the first branch, a softmax function generates dynamic weights for each channel:
$\mathbf{W}_{\text{channel}}=\text{softmax}(\mathbf{C}_{\text{decoded}})$ (5)
In the second branch, $\mathbf{C}_{\text{decoded}}$ is further processed by another 1×1×D sparse convolution:
$\mathbf{C}_{\text{conv}}=\text{SConv}_{1\times 1\times D}(\mathbf{C}_{\text{decoded}})$ (6)
The $\mathbf{W}_{\text{channel}}$ are element-wise multiplied with the output $\mathbf{C}_{\text{conv}}$ to produce the final channel-attended feature map:
$\mathbf{X}_{\text{channel}}=\mathbf{W}_{\text{channel}}\odot\mathbf{C}_{\text{conv}}$ (7)
The final output of the MSCA is obtained by combining the multi-scale fusion output and the channel attention output:
$\mathbf{X}_{\text{output}}=\mathbf{X}_{\text{ms}}+\mathbf{X}_{\text{channel}}$ (8)
@@ III-D Loss Group
Here, $\gamma^{i}$ is decomposed into a class-agnostic parameter $\gamma_{a}$ and a class-specific parameter $\gamma_{b}^{i}$. The parameter $\gamma_{a}$ represents the basic focusing factor under balanced data scenarios, while $\gamma_{b}^{i}\geq 0$ is a variable parameter related to the imbalance degree of class $i$. The term $\eta^{i}=N_{i}/N$ where $N$ is the total number of points in the point cloud, and $N_{i}$ is the number of points in class $i$. The value of $\eta^{i}$ is constrained to the range $[0,1]$, and $1-\eta^{i}$ inversely reflects the weight for low-frequency classes. The hyperparameter $s$ is a scaling factor that determines the upper limit of $\gamma^{i}$.
@@ IV-B Quantitative Results of Curb Detection
In the NRS dataset experiments (refer to Table II), this study compared classic segmentation algorithms such as PointPillars, U-Net, Swin-Transformer, and CSWin-Transformer, as well as the state-of-the-art curb detection model, LCDeT. Leveraging the specially designed MSCA module for 3D curb scenarios, CurbNet achieved the highest detection performance on the NRS dataset. With the auxiliary training of relevant labels, CurbNet attained Precision, Recall, and F-1 scores of 0.8281, 0.8329, and 0.8308, respectively. Among them, auxiliary training helps improve Precision by 0.5 points, and post-processing helps improve Precision by 1.4 points.
@@
The experiments in the 3D-Curb dataset not only compared classic and advanced deep learning model algorithms but also included three traditional methods of manual feature extraction, as shown in Table III. Owing to the robust automatic feature extraction capabilities, deep learning methods significantly outperformed in curb detection, surpassing the other methods by over 10 points in Precision, Recall, and F-1 score. Among the deep learning methods, CurbNet surpassed the best-performing supervised learning model on the SemanticKITTI dataset, Cylinder3D, by 2.5 points in Precision, and exceeded the knowledge distillation model PVKD by 1.5 points. Among them, Precision is improved by more than 1 point through auxiliary training, and by 4.5 points through post-processing.
@@
In our study, we conducted Tolerance performance tests on the 3D-Curb dataset, setting four Tolerances ranging from just 0.05m to 0.2m, as shown in Table IV. Tests were carried out on four models: 3D U-Net, Cylinder3D, PVKD, and CurbNet (ours). With the increase in error Tolerance, performance metrics improved significantly. At 0.05m Tolerance, Precision improved by an average of 4 points; at 0.1m, by 9 points; at 0.15m, by 12 points; and at 0.2m, by 13 points. As Tolerance increased, performance gains gradually reached saturation. Notably, our CurbNet exceeded 0.95 in the average values of Precision, Recall, and F-1 score at just 0.15m Tolerance. This represents the optimal performance achieved in curb detection based on point cloud segmentation.
@@ IV-C Visualization Results of Curb Detection
Furthermore, we conducted a visual analysis of the results on the NRS dataset, as shown in Fig. 8. This analysis primarily showcases curb detection in five common road scenarios and three special intersection types. Through a comparative visualization of the detection results and ground truth, our method accurately identified the respective curb features. The NRS dataset includes several challenging scenes characterized by irregular road structures and unique curb configurations, such as sharp elevation changes, occlusions, and narrow pathways. As shown in Fig. 8 (f)-(h), our method successfully detected the respective curb features in these biased scenarios. This demonstrates the robustness of CurbNet in handling dataset bias and its ability to generalize effectively to complex and irregular road conditions. The consistent detection results in these scenarios validate the model’s capability to adapt to diverse real-world environments, highlighting its potential for broader applicability in intelligent driving systems.
@@ IV-D Post-Processing Experiment
Figure 9: Parameter adjustment experiments for multiple clustering and fitting. We conduct control variable experiments on the main parameter distance variable Eps and the minimum sample point variable minPts of DBSCAN clustering.
@@ IV-E Ablation Study
As shown in Table V, this study conducted comparative ablation experiments focusing on the main loss functions and crucial module designs of our model. We first conducted individual experiments on the employed $\mathcal{L}_{CE}$ (Cross-Entropy Loss), $\mathcal{L}_{FL}$ (Focal Loss), $\mathcal{L}_{ACE}$ (Adaptive Cross-Entropy Loss), and $\mathcal{L}_{IoU}$ (Intersection over Union Loss). Subsequently, we tested combinations of $\mathcal{L}_{CE}$+$\mathcal{L}_{IoU}$ loss, $\mathcal{L}_{FL}$+$\mathcal{L}_{IoU}$ loss and $\mathcal{L}_{ACE}$+$\mathcal{L}_{IoU}$ loss. In the experimental results, due to the design of the $\mathcal{L}_{ACE}$ loss addressing the imbalance in the number of curb point clouds compared to other categories, its disproportionate weight settings caused the model to overly focus on recall during training. However, the interaction with the $\mathcal{L}_{IoU}$ loss led to a more balanced overall performance, achieving optimal detection capabilities. The combination of $\mathcal{L}_{ACE}$+$\mathcal{L}_{IoU}$ loss outperformed the $\mathcal{L}_{CE}$+$\mathcal{L}_{IoU}$ loss group by 1.3 points in Precision and 2.2 points in Recall. It also surpassed the $\mathcal{L}_{FL}$+$\mathcal{L}_{IoU}$ loss group with improvements of 0.7 points in Precision and 2 points in Recall.
Finally, we compared the model’s performance with and without the MSCA module. When the MSCA module was not utilized, the model relied on the original encoder-decoder structure from Cylinder3D [41]. Under the $\mathcal{L}_{ACE}$+$\mathcal{L}_{IoU}$ loss setting, the inclusion of the MSCA module significantly enhanced the performance metrics, particularly increasing Recall by 2 points and the F-1 score by 1 point, which further demonstrates the MSCA module’s effectiveness in improving curb detection performance.